OpenStudy (needhelpinalgebra):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? (5 points) −x − 2y = −4 3x + y = 12

OpenStudy (needhelpinalgebra):

choices: −2y − 4 2y − 4 2y + 4 −2y + 4

OpenStudy (anonymous):

Confusing wording, but it's saying "if you made x= ????? using the first equation, what would be the ?????" -x-2y=-4 Add x to both sides -2y=-4+x add 4 to both sides 4-2y=x

OpenStudy (needhelpinalgebra):

it is saying if you isolated x, making it be by itself. now it wants to use the substitution method, which i can't understand, given the choices.

OpenStudy (needhelpinalgebra):

@AloneS

OpenStudy (anonymous):

I just typed up the answer for you It's -2y+4

OpenStudy (needhelpinalgebra):

oh, cool. Sorry and thanks.